Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the cost of goods completed and transferred out during April
using the weighted average process costing method.
Answer #1

Equivalent units of production = Units transferred + Units in ending WIP x Completion
Material = (18000+80000-44000) + 44000 x 28% = 66320 units
Conversion = (18000+80000-44000) + 44000 x 74% = 86560 units

Cost per unit = Total Cost / Equivalent units
Material = $255332/66320 = $3.85 per unit
Conversion = $385192/86560 = $4.45 per unit

Cost of goods completed and transferred out during April = (18000+80000-44000) x ($3.85+4.45) = $448200